# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ D Derivation of bulk cluster rotation
[…]
@@
Figure S7: Plot fitted angular velocity of the cluster bulk as a function of the cluster radius for three different magnetic field strengths. Solid lines show power law fits.
@@
$\Omega=\frac{4v_{\varphi}(R)\eta}{\Gamma_{v}R^{2}}\propto R^{-2}v_{\varphi}(R).$ (S20)
When edge velocity weakly depends on the cluster radius, as is the case in our system, bulk rotation decays with cluster radius as $R^{-2}$. We demonstrate in Fig. S7 that the fitted angular velocities of our clusters follow this prediction.
E Data evaluation and fitting procedure
@@
$v(x)=v_{\textnormal{max}}e^{-x/\delta}+v_{\textnormal{th}},$ (S21)
with a maximal edge speed $v_{\mathrm{max}}$, decay length $\delta$, and thermal speed $v_{\mathrm{th}}$ being the fitting parameters. To reduce the number of fitting parameters, we followed a two step procedure. (i) We fit the data with Eq. (S21) for several clusters/voids of different size at a fixed field strength. We then compute an average decay length $\delta$ and thermal speed $v_{\mathrm{th}}$ over these clusters/voids. As decay length and thermal speed are assumed to be independent of system size, their average values were taken as true estimates. (ii) Fixing the decay length and thermal speed to their average values, we fit the maximal edge speed $v_{\mathrm{max}}$ again using Eq. (S21). The resulting parameters are shown for different clusters and voids at three different field strengths in Figs. 2, 3 and 4 in the main text.
@@
Here, $\Omega$ is the bulk cluster rotation and $R$ is the effective radius of the cluster determined by fitting the cluster area to a representative circle. The fitting procedure was again split in two steps. (i) We fit the data with Eq. (S22) for several clusters of different size at a fixed field strength. We then compute an average decay length $\delta$ and thermal speed $v_{\mathrm{th}}$. (ii) Fixing these parameters to the average values, we fit the maximal edge speed $v_{\mathrm{max}}$ and the bulk cluster rotation $\Omega$ again using Eq. (S22).
Figure S8: Left: mean speeds in a cluster at $H_{0}=1.1$ mT (black dots) are fitted by two models, without bulk rotation (red line) and with bulk rotation (purple line). Shaded areas indicate the 95% confidence intervals. Right: residuals are plotted for the fit without bulk rotation (top panel) and with bulk rotation (bottom panel). Including bulk rotation makes residuals more symmetric around zero and the sum of their squares smaller.
@@
To confirm that the bulk rotation term is necessary to better describe particle speeds in clusters, we compare two fits, with and without bulk rotation, for a representative cluster. We choose a cluster at an intermediate value of magnetic field $H_{0}=11$ mT and fit Eqs. (S21) and (S22) to its mean speeds in Fig S8. We see that residuals $v_{\mathrm{data}}-v_{\mathrm{fit}}$ are more symmetric around zero, and the sum of their squares is smaller in the fit with bulk rotation than in the fit without rotation. This supports our claim in the main text that the bulk rotation term is necessary to describe macroscopic dynamics in clusters.
F Thermal speed decays with field strength
Figure S9: Thermal speeds $v_{\mathrm{th}}$ obtained from speed fits using Eq. 2 and 4 from the main text for voids (blue dots) and clusters (red dots), respectively. The solid lines indicate linear fits with 95% confidence intervals (shaded).
In this section we demonstrate that thermal speed in both clusters and voids decays with the field strength, as visible in Fig. S9.
G Maximal edge speed and decay length are independent of cluster or void radius
@@
Figure S10: Maximal speed on the edge of a cluster (red) or void (blue) as a function of its radius, predicted from the modified Bessel function solution of the hydrodynamic equation. Parameters are fixed to be $\omega_{p}=1$, $\delta=1$, $\eta/\eta_{R}=50$. Dashed line indicates the maximal speed at flat boundary which is achieved at $R\to\infty$.
@@
Figure S11: Plot fitted maximal speed in clusters and voids as a function of cluster or void radius for three different magnetic field strengths. Solid lines show averages over all clusters or voids of different radii. Shaded regions show the corresponding standard deviation.
In this section we demonstrate that the maximal edge speed and the decay length of edge flows do not depend on the system size for our clusters and sheets with voids.
First, we discuss the case of the maximal edge speed. We consider the solution of the hydrodynamic equations with curved boundary given by Eq. (S17). We plot the maximal velocity in this solution as a function of cluster and void radius in Fig. S10 for all other parameters being fixed. The edge velocity saturates at a constant value in large systems. This value corresponds to a system with flat surface. In the studied systems, the radius is large enough such that the dependence of maximal edge velocity on radius is negligible (Fig. S11). This allows us to treat all clusters and voids as if they have a flat surface.
@@
Figure S12: Plot fitted decay length in clusters and voids as a function of cluster or void radius for three different magnetic field strengths. Solid lines show averages over all clusters or voids of different radii. Shaded regions show the corresponding standard deviation.
@@
We further show that the decay length $\delta$ does not depend on the size of clusters or voids, as we illustrate in Fig. S12. The difference in decay length between clusters and voids, observed in Fig. 4(a) of the main text, happens not due to the difference in their sizes, but because voids are constrained, while the clusters are not.

### v3

@@ Appendix D Derivation of bulk cluster rotation
$\tau_{2}=\int\limits_{0}^{R}dr2\pi r\cdot\Gamma_{v}\Omega r\cdot r=\Omega\Gamma_{v}\pi R^{4}/2,$ (23)
@@
$\Omega=\frac{4v_{\varphi}(R)\eta}{\Gamma_{v}R^{2}}\propto R^{-2}v_{\varphi}(R).$ (24)
When edge velocity weakly depends on the cluster radius, as is the case in our system, bulk rotation decays with cluster radius as $R^{-2}$. We demonstrate in Fig. 7 that the fitted angular velocities of our clusters follow this prediction.
Appendix E Data evaluation and fitting procedure
@@
Figure 9: Thermal speeds $v_{\mathrm{th}}$ obtained from speed fits using Eq. 2 and 4 from the main text for voids (blue dots) and clusters (red dots), respectively. The solid lines indicate linear fits with 95% confidence intervals (shaded).
@@
$v(x)=v_{\textnormal{max}}e^{-x/\delta}+v_{\textnormal{th}},$ (25)
with a maximal edge speed $v_{\mathrm{max}}$, decay length $\delta$, and thermal speed $v_{\mathrm{th}}$ being the fitting parameters. To reduce the number of fitting parameters, we followed a two step procedure. (i) We fit the data with Eq. (25) for several clusters/voids of different size at a fixed field strength. We then compute an average decay length $\delta$ and thermal speed $v_{\mathrm{th}}$ over these clusters/voids. As decay length and thermal speed are assumed to be independent of system size, their average values were taken as true estimates. (ii) Fixing the decay length and thermal speed to their average values, we fit the maximal edge speed $v_{\mathrm{max}}$ again using Eq. (25). The resulting parameters are shown for different clusters and voids at three different field strengths in Figs. 2, 3 and 4 in the main text.
@@
Here, $\Omega$ is the bulk cluster rotation and $R$ is the effective radius of the cluster determined by fitting the cluster area to a representative circle. The fitting procedure was again split in two steps. (i) We fit the data with Eq. (26) for several clusters of different size at a fixed field strength. We then compute an average decay length $\delta$ and thermal speed $v_{\mathrm{th}}$. (ii) Fixing these parameters to the average values, we fit the maximal edge speed $v_{\mathrm{max}}$ and the bulk cluster rotation $\Omega$ again using Eq. (26).
@@
To confirm that the bulk rotation term is necessary to better describe particle speeds in clusters, we compare two fits, with and without bulk rotation, for a representative cluster. We choose a cluster at an intermediate value of magnetic field $H_{0}=11$ mT and fit Eqs. (25) and (26) to its mean speeds in Fig 8. We see that residuals $v_{\mathrm{data}}-v_{\mathrm{fit}}$ are more symmetric around zero, and the sum of their squares is smaller in the fit with bulk rotation than in the fit without rotation. This supports our claim in the main text that the bulk rotation term is necessary to describe macroscopic dynamics in clusters.
@@
Figure 10: Maximal speed on the edge of a cluster (red) or void (blue) as a function of its radius, predicted from the modified Bessel function solution of the hydrodynamic equation. Parameters are fixed to be $\omega_{p}=1$, $\delta=1$, $\eta/\eta_{R}=50$. Dashed line indicates the maximal speed at flat boundary which is achieved at $R\to\infty$.
@@
Figure 11: Plot fitted maximal speed in clusters and voids as a function of cluster or void radius for three different magnetic field strengths. Solid lines show averages over all clusters or voids of different radii. Shaded regions show the corresponding standard deviation.
@@
Figure 12: Plot fitted decay length in clusters and voids as a function of cluster or void radius for three different magnetic field strengths. Solid lines show averages over all clusters or voids of different radii. Shaded regions show the corresponding standard deviation.
Appendix F Thermal speed decays with field strength
In this Appendix, we demonstrate that thermal speed in both clusters and voids decays with the field strength, as visible in Fig. 9.
Appendix G Maximal edge speed and decay length are independent of cluster or void radius
In this Appendix, we demonstrate that the maximal edge speed and the decay length of edge flows do not depend on the system size for our clusters and sheets with voids.
First, we discuss the case of the maximal edge speed. We consider the solution of the hydrodynamic equations with curved boundary given by Eq. (21). We plot the maximal velocity in this solution as a function of cluster and void radius in Fig. 10 for all other parameters being fixed. The edge velocity saturates at a constant value in large systems. This value corresponds to a system with flat surface. In the studied systems, the radius is large enough such that the dependence of maximal edge velocity on radius is negligible (Fig. 11). This allows us to treat all clusters and voids as if they have a flat surface.
@@ Appendix G Maximal edge speed and decay length are independent of cluster or void radius
We further show that the decay length $\delta$ does not depend on the size of clusters or voids, as we illustrate in Fig. 12. The difference in decay length between clusters and voids, observed in Fig. 4(a) of the main text, happens not due to the difference in their sizes, but because voids are constrained, while the clusters are not.